### v1

# Truncated-degree-choosability of planar graphs

Abstract
Assume $G$ is a graph and $k$ is a positive integer. Let $f:V(G)\to\mathbb{N}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$. If $G$ is $f$-choosable, then we say $G$ is $k$-truncated-degree-choosable. It was proved in [Zhou,Zhu,Zhu, Arc-weighted acyclic orientations and variations of degeneracy of graphs, arXiv:2308.15853] that there is a 3-connected non-complete planar graph that is not 7-truncated-degree-choosable, and every 3-connected non-complete planar graph is 16-truncated-degree-choosable. This paper improves the bounds, and proves that there is a 3-connected non-complete planar graph that is not 8-truncated-degree-choosable and every 3-connected non-complete planar graph is 12-truncated-degree-choosable.
Keywords Degree-choosable, Truncated-degree-choosable, Planar graphs
1 Introduction
A list assignment $L$ for a graph $G$ is a mapping that assigns a set $L(v)$ of permissible colours to each vertex $v$ of $G$. We denote by $\mathbb{N}^{G}$ the set of mappings from $V(G)$ to =$\mathbb{N}={\left\{0,1,\ldots\right\}}$. For $f\in\mathbb{N}^{G}$, an $f$-list assignment is an assignment $L$ with $|L(v)|\geq f(v)$ for each vertex $v$. An $L$-colouring of $G$ is a mapping $\phi$ that assigns a colour $\phi(v)$ from $L(v)$ to each vertex $v$, such that $\phi(u)\neq\phi(v)$ for every edge $uv$ of $G$. We say $G$ is $f$-choosable if $G$ is $L$-colourable for any $f$-list assignment $L$ of $G$. The choice number $\operatorname{ch}(G)$ of $G$ is the minimum integer $k$ such that $G$ is $k$-choosable (i.e., $f$-choosable for the constant mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=k$ for all $v$). The choice number is an important graph invariant that has been studied extensively in the literature (cf. [5]).
A graph $G$ is called degree-choosable if it is $f$-choosable for the mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=d_{G}(v)$ for $v\in V(G)$, where $d_{G}(v)$ is the degree of vertex $v$. Degree-choosable graphs are characterized in [3] and [7]: A connected graph $G$ is not degree-choosable if and only if $G$ is a Gallai-tree, i.e., each block of $G$ is either a complete graph or an odd cycle.
In this paper, we investigate a combination of $k$-choosability and degree-choosability, which is called $k$-truncated-degree-choosability (cf. [8]).
Definition 1.
Assume $G$ is a graph and $k$ is a positive integer. Let $f\in\mathbb{N}^{G}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$ for each vertex $v\in V(G)$. If $G$ is $f$-choosable, then we say $G$ is $k$-truncated-degree-choosable.
Truncated-degree-choosability of graphs was first studied by Hutchinson [4]. Bruce Richter asked whether every 3-connected non-complete planar graph is 6-truncated-degree-choosable. Note that for any positive integer $k$, the complete bipartite graph $K_{2,k^{2}}$ is planar and not $k$-truncated-degree-choosable. Also complete graphs are not degree-choosable. So for the study of $k$-truncated choosability of planar graphs, the restriction to 3-connected non-complete graphs is natural. Motivated by Richter’s question, Hutchinson studied truncated-degree-choosability of outerplanar graphs. It was proved in [4] that if $G\neq K_{3}$ is a 2-connected maximal outerplanar graphs (i.e., each inner face is a triangle), then $G$ is 5-truncated-degree-choosable. If $G$ is a 2-connected bipartite outerplanar graphs, then $G$ is 4-truncated-degree-choosable. The results are sharp, as there are 2-connected maximal outerplanar graphs that are not 4-truncated-degree-choosable, and 2-connected bipartite outplanar graphs that are not 3-truncated-degree-choosable. Hutchinson’s result is strengthened in [2], where it is shown that 2-connected outerplanar graphs other than odd cycles are 5-truncated-degree-DP-colourable, and hence 5-truncated-degree-choosable. This result is further strengthened in [6], where it is proved that 2-connected $K_{2,4}$-minor free graphs other than cycles and complete graphs are 5-truncated-degree-DP-colourable. Truncated-degree-choosability of graphs were also studied in [1]. Some families of graphs are shown to be $k$-truncated-degree-choosable, by putting restrictions on the distance between components of the subgraph induced by vertices of degree less than $k$.
Recently, Zhou, Zhu, and Zhu [8] answered Richter’s question in negative and constructed a 3-connected non-complete planar graph that is not 7-truncated-degree-choosable. Then they proved that every 3-connected non-complete planar graph is $ST^{(2)}$-$16$-truncated-degree-degenerate, which implies that these graphs are 16-truncated-degree-choosable.
In this paper, we prove the following results.
Theorem 1.
Every 3-connected non-complete planar graph is 12-truncated-degree-choosable.
Theorem 2.
There is a 3-connected non-complete planar graph which is not 8-truncated-degree-choosable.
We remark that the result in [8] implies that 3-connected non-complete planar graphs are 16-truncated-degree-DP-paintable and 16-truncated-degree-AT. Our result takes advantage of the list assignment, and does not apply to DP-colouring, on-line list colouring and Alon-Tarsi orientation.
2 Some preliminaries
The following lemmas proved in [1] and [8] are needed in our proof. If $G$ is a Gallai-tree, and $L$ is a list assignment of $G$ with $|L(v)|\geq d_{G}(v)$ such that $G$ is not $L$-colourable, then we say $L$ is a bad list assignment for $G$.
Lemma 1 ([1]).
If $G$ is a Gallai-tree, and $L$ is a bad list assignment for $G$, then for each block $B$ of $G$ that is $r$-regular, there is a set $C_{B}$ of $r$ colours such that (i) if $B$ and $B^{\prime}$ share a vertex, then $C_{B}\cap C_{B^{\prime}}=\emptyset$, and (ii) for each vertex $v$, $L(v)=\cup_{v\in B}C_{B}$.
Note that each block of a Gallai-tree is a regular graph. It follows from Lemma 1 that for a bad list assignment $L$ of a Gallai-tree $G$, $|L(v)|=d_{G}(v)$ for each vertex $v\in V(G)$. One property of a bad list assignment $L$ we shall use frequently is that if two vertices $u,v$ are in a same block of $G$ and none of them is a cut-vertex of $G$, then $L(u)=L(v)$.
Assume $G$ is a simple plane graph. We denote by $F(G)$ the set of faces of $G$. For each face $\theta$ of $G$, let $V(\theta)$ be the set of vertices on the boundary of $\theta$. Let $\Theta(G)$ be the bipartite graph with partite sets $V(G)$ and $F(G)$, where $v\theta$ is an edge in $\Theta(G)$ if and only if $v\in V(\theta)$.
Note that $G$ need not be connected. Assume $G$ has connected components $G_{1},G_{2},\ldots,G_{k}$. For a face $\theta$ of $G$, let $J_{\theta}=\{i:V(\theta)\cap V(G_{i})\neq\emptyset\}$. For $i\in J_{\theta}$, let $\theta_{i}$ be the face of $G_{i}$ that contains $\theta$. Then $\theta=\cap_{i\in J_{\theta}}\theta_{i}$ and $V(\theta)=\cup_{i\in J_{\theta}}V\left(\theta_{i}\right)$.
Definition 2.
Let $\theta^{*}$ be the infinite face of $G$, and $v^{*}$ be an arbitrary vertex in $V(\theta^{*})$. A spanning subgraph $F$ of $\Theta(G)$ is very nice with respect to $(\theta^{*},v^{*})$ if the following hold:
  • $d_{F}(v^{*})=1$ and $d_{F}(v)\leq 2$ for $v\in V(G)-\{v^{*}\}$.
  • $d_{F}(\theta^{*})=d_{\Theta(G)}(\theta^{*})$ and $d_{F}(\theta)=d_{\Theta(G)}(\theta)-2$ for each finite face $\theta$ of $G$.
  • If $v_{1},v_{2}\in V(\theta)-N_{F}(\theta)$, then $v_{1},v_{2}$ are contained in a cycle $C$ of $V(\theta)$.
Lemma 2 ([8]).
For any simple plane graph $G$, $\Theta(G)$ has a very nice subgraph.
3 Proof of Theorem 1
We shall prove Theorem 1 by induction on the number of vertices of $G$. For that purpose, we need to conclude that certain induced subgraphs can be properly coloured. However, the induced subgraphs may not be 3-connected. To overcome this difficulty, we shall prove a stronger statement by induction.
Assume $G$ is a connected plane graph, two vertices $x,y$ of $G$ are called visible to each other if $x,y$ lie on the boundary of a same face of $G$, i.e., $G+xy$ is still a planar graph.
Assume $V_{1},V_{2}$ is a partition of $V(G)$ such that each face of $G[V_{2}]$ contains at most one connected component of $G[V_{1}]$. Assume $G_{1},G_{2},\ldots,G_{k}$ are the connected components of $G[V_{2}]$. For a connected component $Q$ of $G[V_{1}]$, let $\theta_{Q}$ be the face of $G[V_{2}]$ that contains $Q$. Let $J_{Q}=\{i:V(\theta_{Q})\cap V(G_{i})\neq\emptyset\}$. For $i\in J_{Q}$, let $\theta_{Q,i}$ be the face of $G_{i}$ which contains $Q$. So $\theta_{Q}=\cap_{i\in J_{Q}}\theta_{Q,i}$. For a face $\theta$ of $G_{i}$, we denote by ${\rm int}(\theta)$ the vertices of $G$ contained in the interior of $\theta$.
Definition 3.
We say $Q$ is properly connected to $V_{2}$ if the following hold:
  1. 1.
    Every vertex in $V({\theta}_{Q})$ is adjacent to some vertex of $Q$.
  2. 2.
    For each $i\in J_{Q}$, if $\theta_{Q,i}$ is a finite face of $G_{i}$, then for each vertex $v$ of $Q$, there are three paths contained in $V(\theta_{Q,i})\cup{\rm int}(\theta_{Q,i})$ connecting $V(\theta_{Q,i})$ and $v$, and these paths are vertex disjoint, except that they have the same end vertex $v$.
Theorem 3.
Assume $G$ is a connected plane graph, and $V_{1},V_{2}$ is a partition of $V(G)$, $V_{2}\neq\emptyset$ and each face of $G[V_{2}]$ contains at most one connected component of $G[V_{1}]$. Assume each connected component $Q$ of $G[V_{1}]$ is properly connected to $V_{2}$. Assume $\theta^{*}$ is the infinite face of $G[V_{2}]$ and $v^{*}\in V(\theta^{*})$. If $f:V(G)\to\mathbb{N}$ is defined as follows:
$f(v)=\begin{cases}1,&\text{ if $v=v^{*}$ and $v^{*}$ is not an isolated vertexin $G[V_{2}]$},\cr 12,&\text{ if $v\in V_{2}-\{v^{*}\}$ or $v=v^{*}$ is an isolated vertex in $G[V_{2}]$},\cr d_{G}(v),&\text{ if $v\in V_{1}$}.\end{cases}$
Then $G$ is $f$-choosable.
Proof of Theorem 1 (using Theorem 3) Assume $G$ is a counterexample to Theorem 1 with minimum number of vertices and subject to this, with maximum number of edges. Let $V_{1}=\{v\in V(G):d_{G}(v)\leq 11\}$ and $V_{2}=V(G)-V_{1}$. Let $L$ be a list assignment of $G$ such that $|L(v)|=d_{G}(v)$ for $v\in V_{1}$ and $|L(v)|=12$ for $v\in V_{2}$, and $G$ is not $L$-colourable. Since $G$ is 3-connected and not a complete graph, $G$ is not a Gallai-tree, and hence $G$ is degree-choosable. Therefore $V_{2}\neq\emptyset$. Also any two vertices in $V_{2}$ visible to each other are adjacent, for otherwise by adding an edge between two non-adjacent vertices in $V_{2}$ that are visible to each other, the resulting graph is not $L$-colourable, and hence is a counterexample with more edges, contrary to the choice of $G$. Therefore each face of $G[V_{2}]$ contains at most one connected component of $G[V_{1}]$, and for each connected component $Q$ of $G[V_{1}]$, each vertex in $V(\theta_{Q})$ is adjacent to some vertex of $Q$.
Assume $Q$ is a connected component $Q$ of $G[V_{1}]$, $i\in J_{Q}$ and $\theta_{Q,i}$ is a finite face of $G_{i}$, and $v$ is a vertex of $Q$. Let $u$ be a vertex of $G$ that lies in the exterior of $\theta_{Q,i}$ (or add such a vertex $u$ adjacent to all vertices in $V(\theta_{Q,i})$ if the exterior of $\theta_{Q,i}$ is empty). Since $G$ is 3-connected, there are three internally vertex disjoint paths connecting $u$ and $v$. Each of these paths has a subpath contained in $V(\theta_{Q,i})\cup{\rm int}(\theta_{Q,i})$ that connects $V(\theta_{Q,i})$ and $v$. These paths are vertex disjoint, except that they have a common end vertex $v$. Hence $Q$ is properly connected to $V_{2}$.
Let $f\in\mathbb{N}^{G}$ be defined as in Theorem 3. As $L$ is an $f$-list assignment of $G$, it follows from Theorem 3 that $G$ is $L$-colourable, a contradiction.
Proof of Theorem 3. Assume Theorem 3 is not true, and $G$ is a counterexample with minimum number of vertices. Let $L$ be an $f$-list assignment of $G$ such that $G$ is not $L$-colourable.
For $X,Y\subseteq V(G)$, let
$N_{Y}(X)=N_{G}(X)\cap Y.$
Definition 4.
Assume $X$ is a subset of $V(G)$ and $\phi$ is an $L$-coloring of $G[X]$. Let $L^{\phi}$ be the list assignment of $G-X$ defined as
$L^{\phi}(v)=L(v)-\{\phi(u):u\in N_{X}(v)\}$
for each vertex $v\in V(G)-X$.
Definition 5.
Assume $X\subseteq Y$ are subsets of $V(G)$, $\phi$ is an $L$-colouring of $G[X]$ and $\psi$ is an $L$-colouring of $G[Y]$. If $\psi(v)=\phi(v)$ for each vertex $v\in X$, then we say $\psi$ is an extension of $\phi$ to $Y$.
Let $F$ be a very nice subgraph of $\Theta(G[V_{2}])$.
Definition 6.
Assume $Q$ is a connected component of $G[V_{1}]$ and $u\in V(\theta_{Q})$. If $u\theta_{Q}$ is an edge of $F$, then we say $u\in V_{2}$ is a protector of $Q$. Otherwise, $u$ is a non-protector of $Q$.
It follows from Lemma 2 that each $u\in V_{2}$ is a protector of at most two connected components of $G[V_{1}]$, $v^{*}$ is a protector of $\theta^{*}$ only, each connected component $Q$ of $G[V_{1}]$ has at most two non-protectors and $\theta^{*}$ has no non-protector. Moreover, if $w_{1},w_{2}$ are two non-protectors of $Q$, then there is a cycle $C_{Q}$ contained in $V(\theta_{Q})$ such that $w_{1},w_{2}\in V(C_{Q})$ and $Q$ is contained in the interior of $C_{Q}$.
Lemma 3.
$G[V_{2}]$ is connected.
  • Proof.
    Assume $G[V_{2}]$ is not connected, and $B$ is a connected component of $G[V_{2}]$ not containing the vertex $v^{*}$. Let $\theta$ be the face of $G[V_{2}]-V(B)$ containing $B$. Let $H$ be the subgraph of $G$ induced by vertices contained in ${\rm int}(\theta)$ and let $G^{\prime}=G-H$. For $i=1,2,$, let $V^{\prime}_{i}=V_{i}\cap V(G^{\prime})$. Then $V^{\prime}_{1},V^{\prime}_{2}$ is a partition of $V(G^{\prime})$ and $V^{\prime}_{2}\neq\emptyset$. Moreover, each connected component of $G^{\prime}[V^{\prime}_{1}]$ is properly connected to $V^{\prime}_{2}$ and $v^{*}\in V(G^{\prime})$. By the minimality of $G$, there is an $L$-colouring $\phi$ of $G^{\prime}$.
    For $i=1,2$, let $V^{\prime\prime}_{i}=V_{i}\cap V(H)$. Then $V^{\prime\prime}_{1},V^{\prime\prime}_{2}$ is a partition of $V(H)$, and $V^{\prime\prime}_{2}\neq\emptyset$. Each connected component of $H[V^{\prime\prime}_{1}]$ is a connected component of $G[V_{1}]$, and is properly connected to $V^{\prime\prime}_{2}$. As
    $|L^{\phi}(v)|\geq\begin{cases}12,&\text{ if $v\in V^{\prime\prime}_{2}$},\cr d_{H}(v),&\text{ if $v\in V^{\prime\prime}_{1}$},\end{cases}$
    by the minimality of $G$, $H$ has an $L^{\phi}$-colouring $\psi$. The union $\phi\cup\psi$ is an $L$-colouring of $G$, a contradiction. ∎
Definition 7.
A leaf block of a connected graph $Q$ is a block $B$ of $Q$ which contains at most one cut-vertex of $Q$. If $B$ contains one cut-vertex $v$ of $Q$, then $v$ is called the root of $B$. The other vertices of $B$ are called non-root vertices. If $Q$ is 2-connected, then $Q$ itself is called a leaf block and all vertices of $Q$ are non-root vertices. For a leaf-block $B$ of $Q$, we denote by $U_{B}$ the set of non-root vertices of $B$.
Lemma 4.
Either $|V_{2}|=1$ or $G[V_{2}]$ is $2$-connected.
  • Proof.
    Assume $|V_{2}|\geq 2$ and $G[V_{2}]$ is not 2-connected. Let $B$ be a leaf-block of $G[V_{2}]$ such that $v^{*}\notin U_{B}$.
    Assume $U_{B}$ is contained in the face $\theta$ of $G[V_{2}]-U_{B}$. If $U_{B}$ is contained in the interior (respectively, exterior) of $\theta$, then let $H$ be the subgraph of $G$ induced by vertices contained in the interior (respectively, exterior) of $\theta$. By the minimality of $G$, there is an $L$-colouring $\phi$ of $G-V(H)$.
    Let $H^{\prime}=H\cup\{v^{\prime}\}$, where $v^{\prime}$ is the root of $B$. For $i=1,2$, let $V^{\prime}_{i}=V_{i}\cap V(H^{\prime})$. Then $V^{\prime}_{1},V^{\prime}_{2}$ is a partition of $V(H^{\prime})$, and $V^{\prime}_{2}\neq\emptyset$. Moreover, $v^{\prime}$ is not an isolated vertex in $H^{\prime}[V^{\prime}_{2}]$. It is easy to see that each connected component of $H^{\prime}[V^{\prime}_{1}]$ is properly connected to $V^{\prime}_{2}$. As
    $|L^{\phi}(v)|\geq\begin{cases}1,&\text{ if $v=v^{\prime}$},\cr 12,&\text{ if $v\in V^{\prime}_{2}$},\cr d_{H}(v),&\text{ if $v\in V^{\prime}_{1}$},\end{cases}$
    by the minimality of $G$, $H$ has an $L^{\phi}$-colouring $\psi$. The union $\phi\cup\psi$ is an $L$-colouring of $G$, a contradiction. ∎
Lemma 5.
Each connected component $Q$ of $G[V_{1}]$ is a Gallai-tree. Moreover, the following hold:
  • Each block of $Q$ is either $K_{n}$ for $n\leq 3$ or is an odd cycle.
  • For each non-cut vertex $v$ of $Q$, $d_{Q}(v)\leq 2$ and $v$ is adjacent to some vertex in $V_{2}$.
  • Proof.
    If a connected component $Q$ of $G[V_{1}]$ is not a Gallai-tree, then by induction hypothesis, $G-Q$ has a proper $L$-colouring $\phi$. As $|L^{\phi}(v)|\geq d_{Q}(v)$ for each vertex $v$ of $Q$ and $Q$ is not a Gallai-tree, then $Q$ admits a proper $L^{\phi}$-colouring $\psi$. The union $\phi\cup\psi$ is a proper $L$-colouring of $G$, a contradiction.
    Assume $Q$ is a connected component of $G[V_{1}]$ and $B$ is a block of $Q$. As $Q$ is a Gallai-tree, $B$ is either a complete graph or an odd cycle. If $B$ is not an odd cycle and not $K_{n}$ for $n\leq 3$, then $B$ is a copy of $K_{4}$. Then $B$ contains a triangle $T$ so that one vertex of $K_{4}$ is contained in the interior of $T$, and all the vertices in the interior of $T$ is not adjacent to $V_{2}$. Let $X$ be the set of vertices contained in the interior of $T$, and let $\phi$ be a proper $L$-colouring of $X$. Then
    $|L^{\phi}(v)|\geq\begin{cases}1,&\text{ if $v=v^{*}$},\cr 12,&\text{ if $v\in V_{2}$},\cr d_{G-X}(v),&\text{ if $v\in V_{1}-X$},\end{cases}$
    It is easy to see that $Q-X$ is properly connected to $V_{2}$. By the minimality of $G$, $G-X$ has a proper $L$-colouring $\psi$ and the union $\phi\cup\psi$ is a proper $L$-colouring of $G$, a contradiction.
    Assume $v\in Q$ is a non-cut vertex of $Q$. Since each block of $Q$ is an odd cycle or a complete graph of order at most $3$, $d_{Q}(v)\leq 2$. Since $Q$ is properly connected to $V_{2}$, $d_{G}(v)\geq 3$. Hence $v$ is adjacent to at least one vertex in $V_{2}$. ∎
Corollary 1.
Assume $Q$ is a connected component of $G[V_{1}]$. If $\theta_{Q}$ is a finite face of $G[V_{2}]$, then $\theta_{Q}=C_{Q}$ is a cycle, and for each leaf block $B$ of $Q$, there are at least two vertices in $C_{Q}$ adjacent to $U_{B}$.
  • Proof.
    As $G[V_{2}]$ is 2-connected, the boundary of each face of $G[V_{2}]$ is a cycle. So $\theta_{Q}=C_{Q}$. Let $B$ be a leaf-block of $Q$ and $v\in U_{B}$. As $Q$ is properly connected to $V_{2}$, there are three paths connecting $C_{Q}$ to $v$, and these paths are pairwise vertex disjoint paths except that they have the same end vertex $v$. At most one of the three paths contains the root vertex of $B$. Each of the other two paths contains an edge between $C_{Q}$ and $U_{B}$. So there are at least two vertices in $C_{Q}$ adjacent to $U_{B}$. ∎
Definition 8.
A subset $X$ of $V(G)$ is valid if for each connected component $Q$ of $G[V_{1}]$, $Q-X$ is connected, and for each vertex $v\in X\cap V(Q)$, $N_{V_{2}}(v)\subseteq X$.
In the following proof process, we shall colour vertices of $G$ one by one, and the set $X$ of coloured vertices will always be a valid subset of $V(G)$. In other word, when we colour a vertex $v$ of a connected component $Q$ of $G[V_{1}]$, and $X$ is the set of already coloured vertices, then $v$ is not a cut-vertex of $Q-X$ and $v$ is not adjacent to any uncoloured vertex of $V_{2}$.
Observation 1.
By following the above rules, if $\phi$ is an $L$-colouring of $G[X]$, then for any connected component $Q$ of $G[V_{1}]$, the following conditions hold:
  1. 1.
    $Q-X$ is a connected component of $G[V_{1}-X]$.
  2. 2.
    For any vertex $v\in V(Q)-X$, $|L^{\phi}(v)|\geq d_{G-X}(v)$.
  3. 3.
    For any vertex $v\in V_{2}-X$, $|L^{\phi}(v)|\geq 12-|N_{V_{2}}(v)\cap X|$.
Definition 9.
Assume $\phi$ is an $L$-colouring of $G[X]$, and $Q$ is a connected component of $G[V_{1}]$. We say $Q$ is free with respect to $\phi$ if for any extension $\psi$ of $\phi$ to $X\cup V_{2}$, $Q-X$ is $L^{\psi}$-colourable.
Given an $L$-colouring $\phi$ of $G[X]$, it is not easy to check whether $Q$ is free with respect to $\phi$. The following are two easy sufficient conditions for $Q$ to be free with respect to $\phi$.
Lemma 6.
Assume $\phi$ is an $L$-colouring of $G[X]$ and $Q$ is a connected component of $G[V_{1}]$. Then $Q$ is free with respect to $\phi$ if one of the following holds:
  1. (P1)
    There is a vertex $v\in V(Q)-X$ for which $|L^{\phi}(v)|>d_{G-X}(v)$.
  2. (P2)
    There are two non-cut vertices $u,v$ of a same block of $Q-X$ with $L^{\phi}(u)\neq L^{\phi}(v)$ and $N_{V_{2}-X}(u)=N_{V_{2}-X}(v)$.
  • Proof.
    By Lemma 5, $Q$ is a Gallai-tree and hence $Q-X$ is also a Gallai-tree.
    If there is a vertex $v\in V(Q)-X$ for which $|L^{\phi}(v)|>d_{G-X}(v)$, then for any extension $\psi$ of $\phi$ to $X\cup V_{2}$, $|L^{\psi}(v)|>d_{Q-X}(v)$, because each time a neighbor of $v$ in $V_{2}$ is coloured, its degree in the subgraph induced by uncoloured vertices decreases by 1, and its list decreases by at most 1. Hence $L^{\psi}|_{Q-X}$ is not a bad list assignment for $Q-X$, and $Q-X$ is $L^{\psi}$-colourable.
Assume there are two non-cut vertices $u,v$ of the same block of $Q-X$ with $L^{\phi}(u)\neq L^{\phi}(v)$ and $N_{V_{2}-X}(u)=N_{V_{2}-X}(v)$. Assume $\psi$ is an extension of $\phi$ to $X\cup V_{2}$. For any $w\in V_{2}$ adjacent to $u$ (hence also adjacent to $v$), either $\psi(w)\in L^{\phi}(u)\cap L^{\phi}(v)$, and or $\psi(w)$ is missing from one of $L^{\phi}(u)$ and $L^{\phi}(v)$. If the later occurs, then one of the following holds:
  • $|L^{\psi}(v)|>d_{Q-X}(v)$.
  • $|L^{\psi}(u)|>d_{Q-X}(u)$.
If the later case never occur, then $L^{\psi}(u)\neq L^{\psi}(v)$. Hence $L^{\psi}|_{Q-X}$ is not a bad list assignment for $Q-X$, and $Q-X$ is $L^{\psi}$-colourable. ∎
Since $G[V_{2}]$ is a planar graph and each planar with more than one vertex has at least two vertices of degree at most $5$, there exists an ordering $<$ of vertices in $V_{2}$ such that $v^{*}$ is the first vertex, and each vertex $u\in V_{2}$ has at most 5 neighbors $v$ with $v<u$
We colour vertices of $G$ one by one and use the following rules to choose the next vertex to be coloured. For $i=1,2,\ldots,$ let $X_{i}$ be the set of the first $i$ coloured vertices and let $\phi_{i}$ be the $L$-colouring of $G[X_{i}]$. Initially, $X_{0}=\emptyset$. Assume $i\geq 0$, $X_{i}$ and $\phi_{i}$ are defined. We construct $X_{i+1}$ and an extension $\phi_{i+1}$ of $\phi_{i}$ to $X_{i+1}$ as follows:
  1. (R1)
    If a non-free connected component $Q$ has a vertex $v\in Q-X_{i}$ which is not a cut-vertex of $Q-X_{i}$, and $v$ is not adjacent to any vertex of $V_{2}-X_{i}$ and $|Q-X_{i}|\geq 2$, then let
    $X_{i+1}=X_{i}\cup\{v\}$
    and let $\phi_{i+1}(v)$ be any colour in $L^{\phi_{i}}(v)$.
  2. (R2)
    If (R1) does not apply, then let $u$ be the smallest vertex in $V_{2}-X_{i}$ with respect to the order $<$, and let
    $X_{i+1}=X_{i}\cup\{u\}.$
For each vertex $u\in V_{2}$, let $i_{u}$ be the index such that
$u\in X_{i_{u}}-X_{i_{u}-1}.$
In other words, $u$ is the $i_{u}$th coloured vertex.
The colour $\phi_{i+1}(u)$ will be selected carefully. The main task of the remaining part of the proof is describing how to choose the colour for $u$ in this case.
Note that when (R1) is applied to colour $v\in Q-X_{i_{u}-1}$, since $|Q-X_{i_{u}-1}|\geq 2$, $|L^{\phi_{i_{u}-1}}(v)|\geq d_{G-X_{i_{u}-1}}(v)\geq 1$. So $L^{\phi_{i_{u}-1}}(v)\neq\emptyset$, and the required colour for $v$ exists. Once a connected component $Q$ of $G[V_{1}]$ becomes free with respect to $\phi_{i}$, then it remains free with respect to $\phi_{j}$ for $j\geq i$. In the following, if the partial colouring $\phi_{i}$ is clear from the context, we simply say $Q$ is free or non-free to mean that $Q$ is free or non-free with respect to $\phi_{i}$.
We apply the above rules until all vertices of $V_{2}$ are coloured. Our goal is to make each connected component $Q$ of $G[V_{1}]$ free when all vertices of $V_{2}$ are coloured.
If this goal is achieved, then $G$ has an $L$-colouring by the definition of free components. To complete the proof of Theorem 3, it remains to show that this goal can be achieved.
Observation 2.
Assume $u\in V_{2}$ and $Q$ is a non-free connected component of $G[V_{1}]$ and $|Q-X_{i_{u}-1}|\geq 2$. Then each non-cut vertex of $Q-X_{i_{u}-1}$ is adjacent to some vertex in $V(\theta_{Q})-X_{i_{u}-1}$.
  • Proof.
    If $v$ is a non-cut vertex of $Q-X_{i_{u}-1}$ which is not adjacent to any vertex in $V(\theta_{Q})-X_{i_{u}-1}$, then (R1) can be applied to colour $v$, in contrary to the definition of $i_{u}$. ∎
Definition 10.
Assume $Q$ is a connected component of $G[V_{1}]$ and $u\in V_{2}$ is a protector of $Q$. Let
$S^{*}_{u,Q}$ $=$ $\{c\in L_{i_{u}-1}(u):\exists\text{ an extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ such that }$
[…]
If $|S^{*}_{u,Q}|\leq 3$, then we say $u$ is a savior of $Q$.
Note that if $Q$ is free with respect to $\phi_{i_{u}-1}$, then $S^{*}_{u,Q}=\emptyset$ and hence $u$ is a savior for $Q$.
To prove that $u$ is a savior for $Q$, it suffices to present a set $S_{u,Q}$ of colours such that $S^{*}_{u,Q}\subseteq S_{u,Q}$ and $|S_{u,Q}|\leq 3$. In other words, the set $S_{u,Q}$ has the following property
Property (S): for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$, $\psi|_{Q-X_{i_{u}-1}}$ is not bad for $Q-X_{i_{u}-1}$.
In the following, for each savior $u$ of a connected component $Q$ of $G[V_{1}]$, a set $S_{u,Q}$ with Property (S) is given. We call $S_{u,Q}$ the colour cost set for $u$ to be a savior of $Q$.
Let
$\mathcal{Q}_{u}=\{Q:\text{ $Q$ is a connected component of $G[V_{1}]$ and $u$ is a savior of $Q$}\}.$
Since $u$ is the protector of at most two connected components $Q$ of $G[V_{1}]$, $|\mathcal{Q}_{u}|\leq 2$.
Now we can finish the description of the colouring process by specifying the colour $\phi_{i_{u}}(u)$ for $u\in V_{2}$: $\phi_{i_{u}}(u)$ is any colour $c$ such that
$c\in L^{\phi_{i_{u}-1}}(u)-\cup_{Q\in\mathcal{Q}_{u}}S_{u,Q}.$
Note that $\cup_{Q\in\mathcal{Q}_{u}}S_{u,Q}$ might be an empty set. In this case, $c$ is any colour in $L^{\phi_{i_{u}-1}}(u)$.
It follows from the colouring rules that $L^{\phi_{i_{u}-1}}(u)=L(u)-\{\phi_{i_{u}-1}(u^{\prime}):u^{\prime}\in V_{2},u^{\prime}<u\}$. Hence
$|L^{\phi_{i_{u}-1}}(u)|=|L(u)|-|\{\phi_{i_{u}-1}(u^{\prime}):u^{\prime}\in V_{2},u^{\prime}<u\}|\geq 12-5=7.$
As $|\mathcal{Q}_{u}|\leq 2$ and $|S_{u,Q}|\leq 3$ for each $Q\in\mathcal{Q}_{u}$, we conclude that
$L^{\phi_{i_{u}-1}}(u)-\cup_{Q\in\mathcal{Q}_{u}}S_{u,Q}\neq\emptyset,$
and hence the required colour $c$ exists.
Lemma 7.
Each connected component $Q$ of $G[V_{1}]$ has a savior.
Assume Lemma 7 is true. We use the colouring strategy described above. When all vertices of $V_{2}$ are coloured, for any connected component $Q$ of $G[V_{1}]$, $Q$ is free. Let $X$ be the set of coloured vertices and $\phi$ be the partial colouring of $X$. Then $Q-X$ has a proper $L^{\phi}$-colouring. Therefore $G$ has a proper $L$-colouring.
So it remains to prove the Lemma 7.
4 Proof of Lemma 7
Assume Lemma 7 is not true and $Q$ is a connected component of $G[V_{1}]$ which has no savior. In particular, for any vertex $u\in V_{2}$, $|L^{\phi_{i_{u}-1}}(v)|=d_{G-X_{i_{u}-1}}(v)$ for each vertex $v\in Q-X_{i_{u}-1}$.
Definition 11.
Assume $u\leq w$ are vertices in $V(\theta_{Q})$ ($u$ and $w$ are not necessarily distinct). We say $w$ is confined by $u$ to colour $c$ if for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ with $\psi(w)\neq c$, $L^{\psi}|_{Q-X_{i_{u}-1}}$ is not bad for $Q-X_{i_{u}-1}$. We may simply say $w$ is confined by $u$, if colour $c$ is clear from the context or is not important.
Since $u\leq w$, at step $i_{u}-1$, $u$ and $w$ are both uncoloured yet. If $w$ is confined by $u$ to colour $c$, then at the step $u$ is to be coloured, we may treat $w$ as coloured with colour $c$ (although $w$ is not coloured yet). Note that if $w$ is confined by $u$ to colour $c$, and $u\leq u^{\prime}\leq w$, then $w$ is confined by $u^{\prime}$ to colour $c$ as well. In particular, if there is a vertex $u$ such that $w$ is confined by $u$ to colour $c$, then $w$ is confined to colour $c$ by itself.
Lemma 8.
Assume $w$ is confined by $u$ to colour $c$. Then the following hold:
  1. 1.
    For any $v\in N_{Q-X_{i_{u}-1}}(w)$, $c\in L^{\phi_{i_{u}-1}}(v)$.
  2. 2.
    $w$ is not a protector of $Q$.
  • Proof.
    (1) If $v\in N_{Q-X_{i_{u}-1}}(w)$ and $c\notin L^{\phi_{i_{u}-1}}(v)$, then for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, either $\psi(w)\neq c$, or $\psi(w)=c\notin L^{\phi_{i_{u}-1}}(v)$ and hence $|L^{\psi}(v)|>d_{Q-X_{i_{u}-1}}(v)$. In any case, $Q-X_{i_{u}-1}$ is $L^{\psi}$-colourable. So $Q$ is free at step $i_{u}-1$, a contradiction.
    (2) If $w$ is a protector of $Q$ which is confined by $u$ to colour $c$, then $w$ is a savior of $Q$ with cost colour set $S_{w,Q}=\{c\}$, a contradiction. ∎
Lemma 9.
Assume $u\leq w\in V(\theta_{Q})$, $v,v^{\prime}$ are non-cut vertices of a same block of $Q-X_{i_{u}-1}$. If $N_{V_{2}-X_{i_{u}-1}}(v)=N_{V_{2}-X_{i_{u}-1}}(v^{\prime})\cup\{w\}$ and $c\in L^{\phi_{i_{u}-1}}(v)-L^{\phi_{i_{u}-1}}(v^{\prime})$, then $w$ is confined by $u$ to colour $c$. Consequently $w$ is not a protector of $Q$.
  • Proof.
    If $\psi$ is an extension of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, and $\psi(w)\neq c$, then either $c\in L^{\psi}(v)-L^{\psi}(v^{\prime})$ and hence $L^{\psi}(v)\neq L^{\psi}(v^{\prime})$, or $v$ and $v^{\prime}$ has a common neighbor $u^{\prime}$ in $V_{2}-X_{i_{u}-1}$ which is coloured by $c\notin L^{\phi_{i_{u}-1}}(v^{\prime})$ and hence $|L^{\psi}(v^{\prime})|>d_{Q-X_{i_{u}-1}}(v^{\prime})$. In any case, it follows from lemma 6 that $Q-X_{i_{u}-1}$ is $L^{\psi}$-colourable. So $w$ is confined by $u$ to colour $c$. By Lemma 8, $w$ is not a protector of $Q$. ∎
Assume $u$ is a protector of $Q$. Let
$F_{u}=\{w\in V_{2}:w\text{ is a non-protector of }Q,w\text{ is confined by }u\}.$
Lemma 10.
Assume $u$ is a protector of $Q$. If $w,w^{\prime}\in F_{u}$, and $w,w^{\prime}$ are confined to a same colour $c$, then $N_{Q-X_{i_{u}-1}}(w)\cap N_{Q-X_{i_{u}-1}}(w^{\prime})=\emptyset$.
  • Proof.
    If $w,w^{\prime}$ are confined to a same colour $c$, and $v\in N_{Q-X_{i_{u}-1}}(w)\cap N_{Q-X_{i_{u}-1}}(w^{\prime})$, then for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, either $\psi(w)\neq c$ or $\psi(w^{\prime})\neq c$ or $|L^{\psi}(v)|>d_{Q-X_{i_{u}-1}}(v)$. In any case, $Q-X_{i_{u}-1}$ is $L^{\psi}$-colourable. So $Q$ is free, a contradiction. ∎
Lemma 11.
Assume $u$ is a protector of $Q$ and $v\in N_{Q-X_{i_{u}-1}}(u)$. Then $d_{G-X_{i_{u}-1}-F_{u}}(v)\geq 4$.
  • Proof.
    Assume to the contrary that $d_{G-X_{i_{u}-1}-F_{u}}(v)\leq 3$. For $w_{i}\in F_{u}$, assume $w_{i}$ is confined by $u$ to colour $c_{i}$. Let
    $S_{u,Q}=L^{\phi_{i_{u}-1}}(v)-\{c_{i}:w_{i}\in F_{u}\text{ and }v\in N_{Q-X_{i_{u}-1}}(w_{i})\}.$
    By Lemma 8, $\{c_{i}:w_{i}\in F_{u},v\in N_{Q-X_{i_{u}-1}}(w_{i})\}\subseteq L^{\phi_{i_{u}-1}}(v)$. By Lemma 10, |$|\{c_{i}:w_{i}\in F_{u},v\in N_{Q-X_{i_{u}-1}}(w_{i})\}|=|\{w_{i}\in F_{u}:v\in N_{Q-X_{i_{u}-1}}(w_{i})\}|$. As $|L^{\phi_{i_{u}-1}}(v)|=d_{G-X_{i_{u}-1}}(v)$, then $|S_{u,Q}|=d_{G-X_{i_{u}-1}-F_{u}}(v)\leq 3$. Let $\psi$ be an extension of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ with $\psi(u)\notin S_{u,Q}$. If $\psi(w_{i})\neq c_{i}$ for some $w_{i}\in F_{u}$, then $Q-X_{i_{u}-1}$ is $L^{\psi}$-colourable.
    If $\psi(w_{i})=c_{i}$ for each $w_{i}\in F_{u}$, then as $\psi(u)\notin S_{u,Q}$, $|L^{\psi}(v)|>d_{Q-X_{i_{u}-1}}(v)$, and hence $L^{\psi}|_{Q-X_{i_{u}-1}}$ is not bad for $Q-X_{i_{u}-1}$. Therefore $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$, a contradiction. ∎
Lemma 12.
$|V(G[V_{2}])|\geq 2$.
  • Proof.
    Assume $V(G[V_{2}])=\{v^{*}\}$, then $|L(v^{*})|=12$ and there is only one face of $G[V_{2}]$ and $G[V_{1}]$ is connected.
    By Lemma 5, each non-cut vertex of $G[V_{1}]$ is adjacent to $v^{*}$. Assume $B$ is a leaf block of $G[V_{1}]$ and $v\in U_{B}$. Then by Lemma 5, $d_{B}(v)\leq 2$ and thus $d_{G}(v)\leq 3$, in contrary to Lemma 11. ∎
Lemma 13.
$\theta_{Q}$ is a finite face of $G[V_{2}]$.
  • Proof.
    Assume $\theta_{Q}$ is the infinite face of $G[V_{2}]$ and $v^{*}\in\theta_{Q}$. Thus each vertex in $\theta_{Q}$ is a protector of $Q$. Let $u$ be the last proctor of $Q$. As $v^{*}$ is the first vertex in the ordering, by Lemma 12, $u\neq v^{*}$ and $v^{*}\in X_{i_{u}-1}$. By Observation 2, each non-cut vertex of $Q-X_{i_{u}-1}$ is adjacent to $u$. Assume $B$ is a leaf block of $Q-X_{i_{u}-1}$ and $v\in U_{B}$. By Lemma 5, $d_{B}(v)\leq 2$ and hence $d_{G-X_{i_{u}-1}}(v)\leq 3$, in contrary to Lemma 11. ∎
By Lemma 13 and Corollary 1, we conclude that $\theta_{Q}=C_{Q}$ and for any block $B$ of $Q$, there are three paths contained in $V(\theta_{Q})\cup V(Q)$ connecting $V(\theta_{Q})$ and $V(B)$, and these paths are pairwise vertex disjoint except that they may have the same end vertex in $B$.
Let $w_{1},w_{2}$ be the non-protectors of $Q$. Since there are three paths connecting $V(\theta_{Q})$ and $Q$, and these paths have the different end vertex in $V(\theta_{Q})$, there is at least one protector of $Q$ adjacent to a vertex in $Q$.
Lemma 14.
Assume $u$ is a protector of $Q$. Then $|Q-X_{i_{u}-1}|\geq 2$.
  • Proof.
    Assume $Q-X_{i_{u}-1}$ consists of a single vertex for some protector $u$ of $Q$. We may assume that $u$ is the last protector of $Q$. Since $Q$ has only two non-protectors, we have $d_{G-X_{i_{u}-1}}(v)=d_{V_{2}-X_{i_{u}-1}}(v)\leq 3$, which contradicts Lemma 11. ∎
Lemma 15.
Assume $u$ is a protector of $Q$. Then each non-cut vertex $v$ of $Q-X_{i_{u}-1}$ is adjacent to at least one non-protector of $Q$.
  • Proof.
    Assume $v$ is a non-cut vertex of $Q-X_{i_{u}-1}$ that is not adjacent to any non-protector of $Q$. By Observation 2, $v$ is adjacent to some vertex of $V_{2}-X_{i_{u}-1}$. Thus $v$ is adjacent to a protector of $Q$. Let $u$ be the last protector of $Q$ which is adjacent to $v$. By Lemma 5, $d_{Q-X_{i_{u}-1}}(v)\leq 2$. Hence $d_{G-X_{i_{u}-1}}(v)\leq 3$, in contrary to Lemma 11. ∎
Lemma 16.
If $B$ is a leaf-block of $Q$, then $w_{1},w_{2}\in N_{\theta_{Q}}(U_{B})$.
  • Proof.
    Assume to the contrary that $w_{2}\notin N_{\theta_{Q}}(U_{B})$. By Corollary 1, $|N_{\theta_{Q}}(U_{B})|\geq 2$. Thus $N_{\theta_{Q}}(U_{B})$ contains a protector of $Q$. Let $u$ be the last protector of $Q$ that is adjacent to some vertex in $U_{B}$.
    By Lemma 11, $d_{G-X_{i_{u}-1}}(v)=|L^{\phi_{i_{u}-1}}(v)|\geq 4$ for each vertex $v\in N_{Q-X_{i_{u}-1}}(u)\cap U_{B}$.
Let $v_{1}\in N_{Q-X_{i_{u}-1}}(u)\cap U_{B}$. As $N_{V_{2}-X_{i_{u}-1}}(v_{1})\subseteq\{u,w_{1}\}$, it follows from Lemma 5, $B-X_{i_{u}-1}$ is an odd cycle $C=[v_{1}v_{2}\ldots v_{2l+1}]$. So $d_{Q-X_{i_{u}-1}}(v_{1})=2$, $d_{G-X_{i_{u}-1}}(v_{1})=4$, $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{1}\}$. As $C$ contains at most one cut-vertex of $Q-X_{i_{u}-1}$, we may assume that $v_{2}\in U_{B}$. By assumption, $N_{V_{2}-X_{i_{u}-1}}(v_{2})\subseteq\{u,w_{1}\}$. By Lemma 15, $v_{2}$ is adjacent to $w_{1}$.
If $v_{2}$ is not adjacent to $u$, then $N_{G-X_{i_{u}-1}}(v_{2})=N_{G-X_{i_{u}-1}}(v_{1})-\{u\}$, and hence $u$ is confined by itself to colour $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$, in contrary to Lemma 9.
Assume $v_{2}$ is also adjacent to $u$. Then $Q-\{v_{1},v_{2}\}$ is contained in the interior of the 4-cycle $[uv_{1}w_{1}v_{2}]$. If $Q$ has another leaf-block $B^{\prime}$, then $|N_{\theta_{Q}}(U_{B^{\prime}})|\leq 1$, a contradiction to Corollary 1. Thus $Q-X_{i_{u}-1}$ is an odd cycle. By Lemma 15, $v_{2l+1}$ is adjacent to $w_{1}$. Thus $v_{2l+1}$ is not adjacent to $u$ (for otherwise $G$ contains $K_{5}$ as a minor). Hence $N_{G-X_{i_{u}-1}}(v_{2l+1})=N_{G-X_{i_{u}-1}}(v_{1})-\{u\}$, and hence $u$ is confined by itself to colour $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2l+1})$, in contrary to Lemma 9. ∎
The following corollary follows from Lemma 16 and the planarity of $G$.
Corollary 2.
$Q$ has at most two leaf-blocks, and each protector $u$ of $Q$ in $V(\theta_{Q})$ is adjacent to non-root vertices of exactly one leaf block of $Q$.
  • Proof.
    If $Q$ has three leaf-blocks, then $Q$ contains $K_{1,3}$ as a minor and by Lemma 16, $Q\cup\{w_{1},w_{2}\}$ contains $K_{3,3}$ as a minor, a contradiction.
Assume $Q$ has two leaf-blocks $B$ and $B^{\prime}$, and $u$ is a protector of $Q$ adjacent to both $U_{B}$ and $U_{B^{\prime}}$. By contracting the cycle $\theta_{Q}$ into a triangle containing $w_{1},w_{2},u$, and contracting each of $U_{B}$ and $U_{B^{\prime}}$ into a single vertex, we obtain a copy of $K_{5}$, a contradiction. ∎
It follows from Corollary 2 that the blocks of $Q$ are ordered as $B_{1},B_{2},\ldots,B_{k}$ ($k\geq 1$) such that $B_{1},B_{k}$ are leaf-blocks and for $i=1,2,\ldots,k-1$, $B_{i}$ and $B_{i+1}$ share a cut-vertex $v_{i}$, and $v_{1},v_{2},\ldots,v_{k-1}$ are pairwise distinct.
In the remainder of this section, let $u$ be the last protector of $Q$.
By Corollary 2, we may assume that $N_{Q}(u)\subseteq U_{B_{k}}$, and assume $v_{1}\in N_{Q}(u)\cap U_{B_{k}}$.
Lemma 17.
$w_{1},w_{2}\notin X_{i_{u}-1}$ and hence $N_{Q}(w_{1})\cup N_{Q}(w_{2})\cap X_{i_{u}-1}=\emptyset$.
  • Proof.
    Assume $w_{2}\in X_{i_{u}-1}$. By Lemma 5, $d_{Q-X_{i_{u}-1}}(v_{1})\leq 2$. By Lemma 11, $d_{G-X_{i_{u}-1}}(v_{1})\geq 4$. It follows that $d_{Q-X_{i_{u}-1}}(v_{1})=2$, $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{1}\}$, $|L^{\phi_{i_{u}-1}}(v_{1})|=d_{G-X_{i_{u}-1}}(v_{1})=4$ and $B_{k}-X_{i_{u}-1}$ is an odd cycle $[v_{1}v_{2}\ldots v_{2l+1}]$.
    Assume first that $U_{B_{k}-X_{i_{u}-1}}-N_{G}(u)\neq\emptyset$, say $v^{\prime}\in U_{B_{k}-X_{i_{u}-1}}-N_{G}(u)$. By Lemma 15, $N_{V_{2}-X_{i_{u}-1}}(v^{\prime})=\{w_{1}\}$ and $|L^{\phi_{i_{u}-1}}(v^{\prime})|=d_{G-X_{i_{u}-1}}(v^{\prime})=3$. Let $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v^{\prime})$. If $\psi$ is an extension of $\phi_{i_{u}-1}$ to $V_{2}\cup X_{i_{u}-1}$ with $\psi(u)\neq c$, then $L^{\psi}(v^{\prime})\neq L^{\psi}(v_{1})$ or $|L^{\psi}(v^{\prime})|>d_{Q-X_{i_{u}-1}}(v^{\prime})$. So $Q-X_{i_{u}-1}$ is $L^{\psi}$-colourable. Therefore $u$ is confined to colour $c$ by itself, a contradiction to Lemma 9.
    Thus $U_{B_{k}-X_{i_{u}-1}}\subseteq N_{G}(u)$. As $u$ has at most two neigbours in $U_{B_{k}-X_{i_{u}-1}}$ (for otherwise $u$ has a neighbour $v_{i}$ with $d_{G-X_{i_{u}-1}}(v_{i})=3$, contrary to Lemma 11), we conclude that $B_{k}-X_{i_{u}-1}$ is a triangle $[v_{1}v_{2}v_{3}]$ and $|U_{B_{k}-X_{i_{u}-1}}|=2$, say $v_{3}$ is a cut-vertex of $G$, and $v_{1},v_{2}\in N_{G}(u)\cap N_{G}(w_{1})$. This implies that $Q$ has at least two blocks, i.e., $k\geq 2$. But then $Q-\{v_{1},v_{2}\}$ is contained in the interior of the 4-cycle $[w_{1}v_{1}uv_{2}]$. Hence $w_{2}$ is not adjacent to $U_{B_{1}}$ in $G$, a contradiction to Lemma 16. ∎
By Lemma 5 and Lemma 15, each non-cut vertex $v$ of $Q$ is adjacent to at least one of $w_{1},w_{2}$. As $w_{1},w_{2}\notin X_{i_{u}-1}$, we have $V(Q)\cap X_{i_{u}-1}=\emptyset$ and $Q-X_{i_{u}-1}=Q$.
Lemma 18.
$Q$ has at least two blocks.
  • Proof.
    Assume to the contrary that $Q$ has a single block.
    If $Q$ is a copy of $K_{2}$ with vertices $v_{1},v_{2}$, then $d_{Q}(v_{i})=1$ and $d_{G-X_{i_{u}-1}}(v_{i})\leq 4$ for $i=1,2$. If $d_{G-X_{i_{u}-1}}(v_{1})=d_{G-X_{i_{u}-1}}(v_{2})=4$, then $\{u,w_{1},w_{2}\}\subseteq N_{V_{2}}(v_{1}),N_{V_{2}}(v_{2})$. Together with $V(\theta_{Q})$, we obtain $K_{5}$ as a minor, a contradiction. Thus we may assume that $d_{G-X_{i_{u}-1}}(v_{2})\leq 3$. Hence $v_{1}$ is adjacent to $u,w_{1},w_{2}$, and by Lemma 11, $v_{2}$ is not adjacent to $u$. If $L^{\phi_{i_{u}-1}}(v_{2})\not\subseteq L^{\phi_{i_{u}-1}}(v_{1})$, then for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, either $|L^{\psi}(v_{1})|>d_{Q}(v_{1})$ or $L^{\psi}(v_{1})\not=L^{\psi}(v_{2})$, and hence $Q$ is $L^{\psi}$-colourable. Thus we may assume that $L^{\phi_{i_{u}-1}}(v_{2})\subseteq L^{\phi_{i_{u}-1}}(v_{1})$. Let
    $S_{u,Q}=L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2}).$
    Then $|S_{u,Q}|\leq 3$. Let $\psi$ be an extension of $\phi_{i_{u}-1}$ to $V_{2}\cup X_{i_{u}-1}$ for which $\psi(u)\notin S_{u,Q}$. We shall show that $Q$ is $L^{\psi}$-colourable, and hence $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
    Assume to the contrary that $L^{\psi}|_{Q}$ is bad for $Q$. Note that $|S_{u,Q}|=|L^{\phi_{i_{u}-1}}(v_{1})|-|L^{\phi_{i_{u}-1}}(v_{2})|=|\{w_{1},w_{2}\}-N_{G}(v_{2})|+1$. If $w_{i}$ is adjacent to $v_{2}$, then $\psi(w_{i})\in L^{\phi_{i_{u}-1}}(v_{2})$, for otherwise $|L^{\psi}(v_{2})|>d_{Q}(v_{2})$. Hence $\psi(w_{i})\notin S_{u,Q}$. So $S_{u,Q}-\{\psi(w_{1}),\psi(w_{2})\}\neq\emptyset$, which implies that $L^{\psi}(v_{1})\neq L^{\psi}(v_{2})$, a contradiction.
Assume $Q$ is an odd cycle $C=[v_{1}v_{2}\ldots v_{2l+1}]$. Note that for each vertex $x\in\{u,w_{1},w_{2}\}$, $N_{Q}(x)$ is a subpath of $C$, and $u$ has at most two neighbours in $C$, for otherwise $u$ has a neighbour $v_{i}$ with $d_{G-X_{i_{u}-1}}(v_{i})=3$, contrary to Lemma 11.
Case 1 $u$ has two neighbours in $C$.
Assume $u$ is adjacent to $v_{1}$ and $v_{2}$. By Lemma 15, for $i\in\{1,2,\ldots,2l+1\}$, $v_{i}$ is adjacent to at least one non-protector. On the other hand, at least one of $v_{1},v_{2}$ is adjacent to only one non-protector, for otherwise $G$ contains a subdivision of $K_{5}$. By symmetry, we assume that $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{u,w_{1}\}$. If $N_{V_{2}-X_{i_{u}-1}}(v_{3})=\{w_{1}\}$, then by Lemma 9, $u$ is confined by itself to colour $c\in L^{\phi_{i_{u}-1}}(v_{2})-L^{\phi_{i_{u}-1}}(v_{3})$, contrary to Lemma 8. If $N_{V_{2}-X_{i_{u}-1}}(v_{3})=\{w_{2}\}$, then $N_{V_{2}-X_{i_{u}-1}}(v_{i})=\{w_{2}\}$ for $i=3,4,\ldots,2l+1$ and $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{2}\}$. Again by Lemma 9, $u$ is confined by itself to colour $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2l+1})$, contrary to Lemma 8. Thus $N_{V_{2}-X_{i_{u}-1}}(v_{3})=\{w_{1},w_{2}\}$. This implies that $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{2}\}$. By symmetry, $N_{V_{2}-X_{i_{u}-1}}(v_{2l+1})=\{w_{1},w_{2}\}$, and this implies that $l=1$, i.e., $C=[v_{1}v_{2}v_{3}]$.
If $L^{\phi_{i_{u}-1}}(v_{1})=L^{\phi_{i_{u}-1}}(v_{2})$, then for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, either $\psi(w_{1})=\psi(w_{2})$ and $|L^{\psi}(v_{3})|>d_{Q}(v_{3})$, or $\psi(w_{1})\neq\psi(w_{2})$, and hence $L^{\psi}(v_{1})\neq L^{\psi}(v_{2})$. In any case, $Q$ is $L^{\psi}$-colourable, and $u$ is a savior for $Q$ with cost colour set $S_{u,Q}=\emptyset$.
If $L^{\phi_{i_{u}-1}}(v_{1})\neq L^{\phi_{i_{u}-1}}(v_{2})$, then let $S_{u,Q}=L^{\phi_{i_{u}-1}}(v_{1})\cap L^{\phi_{i_{u}-1}}(v_{2})$. Then $|S_{u,Q}|\leq 3$, and for any extension $\psi$ of $\phi_{i_{u}-1}$ to $V_{2}\cup X_{i_{u}-1}$ with $\psi(u)\notin S_{u,Q}$, either $|L^{\psi}(v_{1})|>d_{Q}(v_{1})$ or $|L^{\psi}(v_{2})|>d_{Q}(v_{2})$. So $Q$ is $L^{\psi}$-colourable, and $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
Case 2 $u$ has only one neighbour in $C$.
Assume $u$ is adjacent to $v_{1}$. If $v_{1}$ is adjacent to both $w_{1}$ and $w_{2}$, then there exists $i\in\{2,2l+1\}$ such that $d_{G-X_{i_{u}-1}}(v_{i})=3$. By symmetry, we may assume that $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{w_{1}\}$. Let $S_{u,Q}$ be a subset of $L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$ of size $2$. For any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$, at least one of the following holds:
  • $\psi(w_{1})\notin L^{\phi_{i_{u}-1}}(v_{2})$ and hence $|L^{\psi}(v_{2})|>d_{Q}(v_{2})$.
  • $\psi(w_{1})\in L^{\phi_{i_{u}-1}}(v_{2})$ and hence $L^{\psi}(v_{1})\not=L^{\psi}(v_{2})$, because at least one of the two colours in $S_{u,Q}$ is contained in $L^{\psi}(v_{1})-L^{\psi}(v_{2})$.
So $Q$ is $L^{\psi}$-colourable, and $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
Assume $v_{1}$ is adjacent to only one of $w_{1},w_{2}$. By symmetry, assume that $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{1}\}$.
If $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{w_{1}\}$, then it follows from Lemma 9 that $u$ is confined by itself to colour $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$, a contradiction. If $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{w_{1},w_{2}\}$, then $w_{2}$ is adjacent to $v_{2},v_{3},\ldots,v_{2l+1}$, and $N_{V_{2}-X_{i_{u}-1}}(v_{2l+1})=\{w_{2}\}$. It follows from Lemma 9 that $w_{1}$ is confined by $u$ to colour $c\in L^{\phi_{i_{u}-1}}(v_{2})-L^{\phi_{i_{u}-1}}(v_{2l+1})$. But then $d_{G-X_{i_{u}-1}-F_{u}}(v_{1})=3$, contrary to Lemma 11.
So $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{w_{2}\}$. By symmetry, $N_{V_{2}-X_{i_{u}-1}}(v_{2l+1})=\{w_{2}\}$. Hence for $i=2,3,\ldots,2l+1$, $N_{V_{2}-X_{i_{u}-1}}(v_{i})=\{w_{2}\}$.
Assume $w_{1}$ and $w_{2}$ are adjacent. If $L^{\phi_{i_{u}-1}}(v_{2})\subseteq L^{\phi_{i_{u}-1}}(v_{1})$, then let $S_{u,Q}=L^{\phi_{i_{u}-1}}(v_{2})$. Let $\psi$ be an extension of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$. Then since $\psi(w_{1})\neq\psi(w_{2})$, then either ${\psi}(u)\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$ and hence $L^{\psi}(v_{1})\neq L^{\psi}(v_{2})$ or ${\psi}(u)\notin L^{\phi_{i_{u}-1}}(v_{1})$ and hence $|L^{\psi}(v_{1})|>d_{Q}(v_{1})$. In any case, $Q$ is $L^{\psi}$-colourable, and $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
If $L^{\phi_{i_{u}-1}}(v_{2})\not\subseteq L^{\phi_{i_{u}-1}}(v_{1})$, then $|L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})|\geq 2$. Let $S_{u,Q}$ be a subset of $L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$ of size $2$. For any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$, at least one of the colours in $S_{u,Q}$ is contained in $L^{\psi}(v_{1})-L^{\psi}(v_{2})$. Hence $L^{\psi}(v_{1})\neq L^{\psi}(v_{2})$ and $Q$ is $L^{\psi}$-colourable. So $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
Now we assume $w_{1}$ and $w_{2}$ are not adjacent. Then $|\theta_{Q}|\geq 4$, and $\theta_{Q}$ contains another protector of $Q$. As there are three pairwise vertex disjoint paths connecting $\theta_{Q}$ to $v_{2}$ except that have the same end vertex $v_{2}$, at most one of the three paths contains $w_{2}$ and at most one of the three paths contains $v_{1}$. Thus $Q$ has a protector which is adjacent to $v_{2}$. By symmetry, $Q$ has a protector which is adjacent to $v_{2l+1}$. Let $w$ be the last protector of $Q$ which is adjacent to $v_{2}$ or $v_{2l+1}$. If $w$ is adjacent to both $v_{2}$ and $v_{2l+1}$, then $V(Q)\cup V(\theta_{Q})$ contains $K_{5}$ as a minor, a contradiction. By symmetry, we assume that $w$ is adjacent to $v_{2}$. Thus
$N_{V_{2}-X_{i_{w}-1}}(v_{2l+1})=N_{V_{2}-X_{i_{w}-1}}(v_{2})-\{w\}.$
This contradicts Lemma 9. ∎
It follows from Corollary 2 and Lemma 18 that $k\geq 2$ and $Q$ has exactly two leaf-blocks $B_{1}$ and $B_{k}$. Recall that $v_{1}\in N_{Q}(u)\cap U_{B_{k}}$ and $N_{Q}(u)\subseteq U_{B_{k}}$. We complete the proof of Theorem 3 by considering two cases.
Case 1 $|L^{\phi_{i_{u}-1}}(v_{1})|=d_{G-X_{i_{u-1}}}(v_{1})=5$.
Then $B_{k}$ is an odd cycle $C=[v_{1}v_{2}\ldots v_{2l+1}]$ and $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{1},w_{2}\}$. By symmetry, we may assume that $v_{2}\in U_{B_{k}}$.
By the planarity of $G$, $v_{2}$ is adjacent to only one of $u,w_{1},w_{2}$. By Lemma 11, $v_{2}$ is not adjacent to $u$. By symmetry, assume $N_{V_{2}-X_{i_{u}-1}}(v_{2})=\{w_{1}\}$. Then $|L^{\phi_{i_{u}-1}}(v_{2})|=d_{G-X_{i_{u}-1}}(v_{2})=3.$ Let $S_{u,Q}$ be a subset of $L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v_{2})$ of size $2$. For any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$, at least one of the following holds:
  • $\psi(w_{1})\notin L^{\phi_{i_{u}-1}}(v_{2})$ and hence $|L^{\psi}(v_{2})|>d_{Q}(v_{2})$.
  • $\psi(w_{1})\in L^{\phi_{i_{u}-1}}(v_{2})$ and hence $L^{\psi}(v_{1})\not=L^{\psi}(v_{2})$.
So $Q$ is $L^{\psi}$-colourable, and $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
Case 2 $|L^{\phi_{i_{u}-1}}(v_{1})|=d_{G-X_{i_{u-1}}}(v_{1})=4$.
First we show that $B_{1}$ has a non-root vertex $v^{\prime}_{1}$ with $d_{G-X_{i_{u}-1}}(v^{\prime}_{1})=3$.
If $B_{1}=K_{2}$, then $B_{1}$ has a single non-root vertex $v^{\prime}_{1}$. By Lemma 16, $v^{\prime}_{1}$ is adjacent to both $w_{1}$ and $w_{2}$, and not adjacent to $u$. Hence $d_{G-X_{i_{u}-1}}(v^{\prime}_{1})=3.$
If $B_{1}$ is an odd cycle $[v^{\prime}_{1}v^{\prime}_{2}\ldots v^{\prime}_{2t+1}]$, then we may assume that $v^{\prime}_{1},v^{\prime}_{2}$ are non-root vertices of $B_{1}$, and at most one of $v^{\prime}_{1},v^{\prime}_{2}$ is adjacent to both $w_{1},w_{2}$ (for otherwise $V(\theta_{Q})\cup V(Q)$ contains $K_{5}$ as a minor, a contradiction). Thus we may assume that $|L^{\phi_{i_{u}-1}}(v^{\prime}_{1})|=d_{G-X_{i_{u}-1}}(v^{\prime}_{1})=3$.
Case 2(i) For all the other non-root vertex $v^{\prime}$ of $B_{1}$, $L^{\phi_{i_{u}-1}}(v^{\prime})=L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$.
Let $c\in L^{\phi_{i_{u}-1}}(v_{1})-L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$.
Let $l$ be the largest index such that $B_{l}$ has a vertex $x$ such that $c\notin L^{\phi_{i_{u}-1}}(x)$. Since $c\notin L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$, we know that $l\geq 1$ is well-defined. Let
$S_{u,Q}=\begin{cases}\{c\},&\text{ if $k-l$ is even},\cr L^{\phi_{i_{u}-1}}(v_{1})-\{c\},&\text{ if $k-l$ is odd}.\end{cases}$
Assume $\psi$ is an extension of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ such that $\psi(u)\notin S_{u,Q}$, and $L^{\psi}|_{Q}$ is a bad list assignment for $Q$. By Lemma 1, for $1\leq i\leq q$, there is a set $C_{i}$ of colours such that for each vertex $x$ of $Q$, $L^{\psi}(x)=\cup_{x\in B_{i}}C_{i}$ and $C_{i}\cap C_{i+1}=\emptyset$.
For $i=1,2$, we have $\psi(w_{i})\in L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$, for otherwise since $L^{\phi_{i_{u}-1}}(v^{\prime})=L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$ for all non-root vertex $v^{\prime}$ of $B_{1}$, and each $w_{i}$ is adjacent to some non-root vertex of $B_{1}$, we conclude that there is a non-root vertex $v^{\prime}$ of $B_{1}$ such that $|L^{\psi}(v^{\prime})|>d_{Q}(v^{\prime})$, in contrary to the assumption that $L^{\psi}|_{Q}$ is a bad list assignment for $Q$. So $\psi(w_{i})\neq c$ for $i=1,2$.
Since $w_{1},w_{2}\in N_{G}(U_{B_{i}})$ for $i=1,k$, the subgraph of $G$ induced by $\{w_{1},w_{2}\}\cup U_{B_{1}}\cup U_{B_{q}}$ contains a cycle $C^{\prime\prime}$ such that $Q-U_{B^{\prime}_{1}}-U_{B^{\prime}_{q}}$ are contained in the interior of $C^{\prime\prime}$, and $u$ is contained in the exterior of $C^{\prime\prime}$. Thus $u$ is not adjacent to any vertex in $Q-U_{B^{\prime}_{1}}-U_{B^{\prime}_{q}}$. So for any vertex $x$ of $Q-U_{B^{\prime}_{1}}-U_{B^{\prime}_{q}}$,
$L^{\phi_{i_{u}-1}}(x)-\{\psi(w_{1}),\psi(w_{2})\}\subseteq L^{\psi}(x).$
In particular, $c\in L^{\psi}(x)$ if and only if $c\in L^{\phi_{i_{u}-1}}(x)$.
As there is a vertex $x\in B_{l}$ such that $c\notin L^{\phi_{i_{u}-1}}(x)$, we know that $c\notin C_{l}$.
As $c\in L^{\phi_{i_{u}-1}}(x)$ for all $x\in B_{l+1}$, we know that $c\in L^{\psi}(x)$ for all $x\in B_{l+1}$. This implies that $c\in C_{l+1}$. As $C_{l+1}\cap C_{l+2}=\emptyset$, we know that $c\notin C_{l+2}$. Now $c\in L^{\psi}(x)$ for all $x\in B_{l+3}$. This implies that $c\in C_{l+3}$. Repeat this argument, we conclude that $c\in C_{l+1},C_{l+3},\ldots,C_{l+1+2t},\ldots$. If $k-l$ is odd, this implies that $c\in C_{k}$, and hence $\psi(u)\neq c$. Therefore $\psi(u)\notin L^{\phi_{i_{u}-1}}(v_{1})$ and $|L^{\psi}(v_{1})|>d_{Q}(v_{1})$. So $L^{\psi}|_{Q}$ is not a bad list assignment for $Q$, a contradiction.
If $k-l$ is even, then $c\notin C_{l},C_{l+2},\ldots,C_{k}$. But $\psi(u)\notin S_{u,Q}=\{c\},\psi(w_{i})\neq c$ for $i=1,2$. Hence $c\in L^{\psi}(v_{1})$, a contradiction.
Case 2(ii) $L^{\phi_{i_{u}-1}}(v^{\prime}_{1})\neq L^{\phi_{i_{u}-1}}(v^{\prime}_{2})$ for $v^{\prime}_{1},v^{\prime}_{2}\in U_{B_{1}}$.
In this case $B_{1}$ is an odd cycle $C^{\prime}=[v^{\prime}_{1}v^{\prime}_{2}\ldots v^{\prime}_{2t+1}]$. Note that $N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{1})\neq N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{2})$, for otherwise for any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$, we have either $L^{\psi}(v^{\prime}_{1})\neq L^{\psi}(v^{\prime}_{2})$ or $|L^{\psi}(v^{\prime}_{i})|>d_{Q}(v^{\prime}_{i})$ for some $i\in\{1,2\}$, and hence $u$ is a savior of $Q$ with cost colour set $S_{u,Q}=\emptyset$.
If one of the $v^{\prime}_{1},v^{\prime}_{2}$ is adjacent to both $w_{1}$ and $w_{2}$, without loss of generality, we assume $N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{2})=\{w_{1},w_{2}\}$ and $N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{1})=\{w_{2}\}$. By Lemma 9, $w_{1}$ is confined by $u$ to colour $c\in L^{\phi_{i_{u}-1}}(v^{\prime}_{2})-L^{\phi_{i_{u}-1}}(v^{\prime}_{1})$.
If $B_{k}$ has a non-root vertex $v$ which is adjacent to $u$ and $w_{1}$, then $d_{G-X_{i_{u}-1}-F_{u}}(v)=3$, contrary to Lemma 11.
If each non-root vertex of $B_{k}$ is not adjacent to both $u$ and $w_{1}$, then $B_{k}$ is an odd cycle $[v_{1}v_{2}\ldots v_{2l+1}]$. Since $v_{1}\in N_{Q}(u)\cap U_{B_{k}}$, we assume that $N_{V_{2}-X_{i_{u}-1}}(v_{1})=\{u,w_{2}\}$. Since $w_{1},w_{2}\in N_{V_{2}}(U_{B_{k}})$, there exists a vertex $v_{i}\in U_{B_{k}}$ such that $N_{V_{2}-X_{i_{u}-1}}(v_{i})=\{w_{1}\}$. It follows from Lemma 8 that $c\in L^{\phi_{i_{u}-1}}(v_{i})$. Then let $S_{u,Q}$ be a subset of $L^{\phi_{i_{u}-1}}(v_{1})-(L^{\phi_{i_{u}-1}}(v_{i})-\{c\})$ of size $2$.
For any extension $\psi$ of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{2}$ for which $\psi(u)\notin S_{u,Q}$, $\psi(w_{1})\not=c$ or $\psi(w_{1})=c$ but then at least one of the colours in $S_{u,Q}$ is contained in $L^{\psi}(v_{1})-L^{\psi}(v_{i})$ and hence $L^{\psi}(v_{1})\neq L^{\psi}(v_{i})$. In any case $Q$ is $L^{\psi}$-colourable. So $u$ is a savior for $Q$ with cost colour set $S_{u,Q}$.
Assume each of $v^{\prime}_{1}$ and $v^{\prime}_{2}$ is adjacent to exactly one non-protector, say $N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{1})=\{w_{2}\}$ and $N_{V_{2}-X_{i_{u}-1}}(v^{\prime}_{2})=\{w_{1}\}$. Assume $\psi$ is an extension of $\phi_{i_{u}-1}$ to $X_{i_{u}-1}\cup V_{i}$ such that $L^{\psi}|_{Q}$ is bad for $Q$. As $L^{\phi_{i_{u}-1}}(v^{\prime}_{1})\neq L^{\phi_{i_{u}-1}}(v^{\prime}_{2})$ and $L^{\psi}(v^{\prime}_{1})=L^{\psi}(v^{\prime}_{2})$, we conclude that $w_{1}$ is confined by $u$ to the unique colour $c_{1}\in L^{\psi}(v^{\prime}_{1})-L^{\psi}(v^{\prime}_{2})$ and $w_{2}$ is confined by $u$ to the unique colour $c_{2}\in L^{\psi}(v^{\prime}_{2})-L^{\psi}(v^{\prime}_{1})$. Thus $w_{1},w_{2}\in F_{u}$, and hence $d_{G-X_{i_{u}-1}-F_{u}}(v_{1})\leq 3$, contrary to Lemma 11. This completes the proof of Lemma 7. ∎
5 Proof of Theorem 2
In this section, we present a 3-connected non-complete planar graph which is not 8-truncated-degree-choosable.
Fig. 1: The graph $H$
Let $H$ be the graph in Figure 1 and let $L$ be the list assignment of $H$ defined as follows:
  • $L(x)=\{a\}$, $L(y)=\{b\}$,
  • $L(u_{i})=L(v_{i})=\{a,b,1,2,3,4,5,6\}$ for $i=1,2$,
  • $L(s_{1})=L(t_{1})=\{a,1,2,3\}$, $L(w_{1})=\{a,4,5,6\}$,
  • $L(s_{2})=L(t_{2})=L(s_{3})=L(t_{3})=\{1,2,3\}$, $L(w_{2})=L(s_{6})=L(t_{6})=\{4,5,6\}$,
  • $L(s_{4})=L(t_{4})=\{b,1,2,3\}$, $L(s_{7})=L(t_{7})=\{b,4,5,6\}$,
  • $L(s_{5})=L(t_{5})=\{b,1,2,3,7\}$, $L(s_{8})=L(t_{8})=\{b,4,5,6,7\}$.
First we show that $H$ is not $L$-colourable. Assume to the contrary that there is an $L$-colouring $\phi$ of $H$.
Claim 1.
$\phi(u_{1})\in\{1,2,3\}$ or $\phi(v_{1})\in\{1,2,3\}$.
  • Proof.
    Let $H_{1}=H[\{u_{1},v_{1},w_{1},w_{2}\}]$. Then $\phi$ is an $L_{1}$-colouring of $H_{1}$ where $L_{1}(u_{1})=L_{1}(v_{1})=\{1,2,3,4,5,6\}$ and $L_{1}(w_{1})=L_{1}(w_{2})=\{4,5,6\}$. Assume $\phi(u_{1})\notin\{1,2,3\}$ and $\phi(v_{1})\notin\{1,2,3\}$. Then $\{\phi(u_{1}),\phi(v_{1}),\phi(w_{1}),\phi(w_{2})\}\subseteq\{4,5,6\}$. But $\{u_{1},v_{1},w_{1},w_{2}\}$ induces a copy of $K_{4}$, a contradiction. ∎
By symmetry, we assume that $\phi(u_{1})\in\{1,2,3\}$.
Observation 3.
Assume $K$ is a copy of $K_{4}$ with vertex set $\{x_{1},x_{2},x_{3},x_{4}\}$, and $L^{\prime}$ is a list-assignment of $K$ such that $L^{\prime}(x_{1})=L^{\prime}(x_{2})=L^{\prime}(x_{3})=\{a_{1},a_{2},a_{3}\}$ and $L^{\prime}(x_{4})=\{a_{1},a_{2},a_{3}\}\cup B$ (where $B$ is a non-empty colour set and $B\cap\{a_{1},a_{2},a_{3}\}=\emptyset$). Then in each $L^{\prime}$-colouring $\varphi$ of $K$, $\varphi(x_{4})\in B$.
Let $H_{2}=H[u_{1},u_{2},s_{1},s_{2}]$. Then $\phi$ is an $L_{2}$-colouring of $H_{2}$ where $L_{2}(u_{1})=L_{2}(s_{1})=L_{2}(s_{2})=\{1,2,3\}$ and $L_{2}(u_{2})=\{1,2,3,4,5,6\}$. By Observation 3, $\phi(u_{2})\in\{4,5,6\}$.
Let $H_{3}=H[u_{1},s_{3},s_{4},s_{5}]$. Then $\phi$ is an $L_{3}$-colouring of $H_{3}$ where $L_{3}(u_{1})=L_{3}(s_{3})=L_{3}(s_{4})=\{1,2,3\}$ and $L_{3}(s_{5})=\{1,2,3,7\}$. By Observation 3, $\phi(s_{5})=7$.
Similarly, let $H_{4}=H[u_{2},s_{6},s_{7},s_{8}]$. Then $\phi$ is an $L_{4}$-colouring of $H_{4}$ where $L_{4}(u_{2})=L_{4}(s_{6})=L_{4}(s_{7})=\{4,5,6\}$ and $L_{4}(s_{8})=\{4,5,6,7\}$. By Observation 3, $\phi(s_{8})=7$.
Thus $\phi(s_{5})=\phi(s_{8})=7$, but $s_{5}s_{8}$ is an edge in $H$, a contradiction. Hence $H$ is not $L$-colourable.
Let $G$ be a graph obtained from the disjoint union of 56 copies of $H_{i}$ of $H$ by identifying all copies of $x$ into a single vertex (also named as $x$) and all the copies of $y$ into a single vertex (also named as $y$), and adding the edges $v_{2}^{(i)}u_{2}^{(i+1)}$ (where $v_{2}^{(i)}$ and $u_{2}^{(i)}$ are the copies of $u_{2}$ and $v_{2}$ in $H_{i}$) for $i=1,2,\dots,55$, and adding an edge connecting $x$ and $y$. Then $G$ is a non-complete $3$-connected planar graph.
Let $L(x)=L(y)=\{a,b,c,d,e,f,g,h\}$. There are 56 possible $L$-colourings $\phi$ of $x$ and $y$. Each such a colouring $\phi$ corresponds to one copy of $H$. We define the list assinment of the corresponding copy of $H$ as $L$ by replacing $a$ with $\phi(x)$ and replacing $b$ with $\phi(y)$. It is easy to verify that $L(v)=\min\{d(v),8\}$ for any $v\in V(G)$. As every possible $L$-colouring of $x$ and $y$ cannot be extended to an $L$-colouring of some copy of $H$, we conclude that $G$ is not $L$-colourable. Hence $G$ is not $8$-truncated-degree-choosable.